21) A point-like charged object with Q1 = +6.0 nC is at the point (0.30 m, 0.00 m); an object with charge Q2 = -1.0 nC is at (0.00 m, 0.10 m), and an object with charge Q3 = +5.0 nC is at (0.00 m, 0.00 m). What are the magnitude and direction of the net force on the +5.0-nC object due to the other two objects? (k = 9.0 × 109 N∙m2/C2)
Answer #1

force due to q1

F1 = k q1 q3 / r1^2

F1 = 9*10^9* 6*10^-9* 5*10^-9 / ( 0.3^2)

F1 = 3*10^-6 N

force due to q2

F2 = k q2 q3 / r2^2

F2 = 9*10^9* 5*10^-9* 1*10^-9 / ( 0.1^2)

F2 = 4.5*10^-6 N

net force is given by

F^2 = F1^2 + F2^2

F = 5.408*10^-6 N ( Ans)

========

Direction,

phi = 180 - arctan ( F2 / F1)

phi = 123.7 ccw from +x axis

=======

Comment in case any doubt, will reply for sure . Goodluck